A shooter aims horizontally at a target at the dis
Answer #1

Horizontal distance = d
Initial Speed of bullet, = v

So time taken to travel distance, t = Distance/Speed = d/v

Now,
Initial Vertical Speed = 0
Vertical acceleration = g = 9.8 m/s^2

So Height by which target will be missed,
s = u*t + 1/2*a*t^2
h = 0 + 1/2 g * (d/v)^2

h = 1/2*g * (d/v)^2 or 4.9 * (d/v)^2